# EgoEMS: A High-Fidelity Multimodal Egocentric Dataset for Cognitive Assistance in Emergency Medical Services

### v1

@@ D.2 Keystep Segmentation
Model (Architecture) Modality (Feature Extraction) IOU Evaluation Metric
@@
MTRSAP (Transformer) weerasinghe2024multimodal Ego Video (Resnet50) 10% 0.54 0.53 0.53 0.55 0.55
@@
VideoLlama-3.3 (VLM) zhang2025videollama Ego Video (End-to-End) 10% 0.07 0.51 0.12 0.33 0.34
@@ D.2.3 Audio-Based Zero-Shot Setting
Method. As shown in the Figure 33, our method contains three main modules: 1) Whisper-timestamped lintoai2023whispertimestamped takes audio $x_{\text{window}}$ as input and conducts automatic speech recognition with word-level timestamps; 2) GPT-4o achiam2023gpt is prompted in Chain-of-Thought wei2022chain to decompose keystep detection task into two sub-tasks ge-etal-2024-dkec: token classification and phrase span detection and ask LLMs to think step by step. The input for GPT-4o is the transcripts generated by Whisper-timestamped, and the output is the phrases from the transcripts, each labeled with keystep type. The detailed prompt is shown in Figure 34. 3) Keystep Boundary Detection further applies keystep span detection by grouping phrases with the same keystep label together and updating the start and end timestamp for each grouped keystep. The predictions for all segments $x_{\text{window}}$ are concatenated to form a complete trial, which is then passed to the final Keystep Boundary Detection module.
@@ D.3 CPR quality estimation
Figure 35: Multimodal fusion approach for CPR quality estimation.
Rate (cpm) Depth (mm) F1 Acc F1 Acc
Egocentric (RGB) Heuristic (SP) 28.56 25.78 0.47 0.52 0.77 0.70
IMU (Acc) Supervised + SP 14.89 12.13 0.51 0.64 0.83 0.77
Fusion (Supervised) Supervised + SP 15.66 11.78 0.52 0.62 0.83 0.77
@@ D.3.1 Egocentric Setting
Supervised depth estimation. For this approach, we employ a supervised depth estimation model to estimate the depth of RGB frames. Specifically, we utilize a monocular depth estimation method lasinger2019towards to obtain a relative depth map, which we scale to absolute depth using a reference object with known real-world depth. Further, we evaluate the large and small variants of the MiDas model introduced in lasinger2019towards.
@@ D.3.2 Smartwatch IMU-Based Setting
Figures 38 and 39 present the average errors in compression rate and depth estimation for different modalities, grouped by participant background. In contrast to egocentric video based estimation, smartwatch IMU based method consistently achieve low errors for both EMS responders and general public participants. This improved performance stems from the direct and robust measurement of wrist motion by the IMU, which is largely unaffected by the occlusions, field-of-view limitations, and other challenges that hinder egocentric video based approaches. This is further evident from the substantial reduction in compression rate error for EMS responders when using IMU data compared to egocentric video. Following a similar trend, the smartwatch IMU based method achieves the highest feedback accuracy for compression rate at 0.64, and for compression depth at 0.77, outperforming other modalities in both metrics.
@@ D.3.3 Multimodal Fusion
These quantitative results are further supported by the box plots in Figures 38 and 39, which illustrate deviations in predicted CPR compression rate and depth across modalities and participant backgrounds. For both metrics, fusion methods exhibit consistently lower deviations compared to unimodal egocentric video, especially for EMS responders, whose results with ego video alone show large variability and frequent extreme errors. The addition of IMU data stabilizes predictions, reducing errors caused by wrist occlusions, camera motion, and confusion with secondary responders in multi-responder scenes.

### v2

@@ D.2 Keystep Segmentation
Learning Type Model (Architecture) Modality (Feature Extraction) IOU Evaluation Metric
@@
Supervised MTRSAP (Transformer) weerasinghe2024multimodal Ego Video (Resnet50) 10% 0.54 0.53 0.53 0.55 0.55
@@
Zero-shot VideoLlama-3.3 (VLM) zhang2025videollama Ego Video (End-to-End) 10% 0.07 0.51 0.12 0.33 0.34
@@ D.2.3 Audio-Based Zero-Shot Setting
Method. As shown in the Figure 33, our method contains three main modules: i) Whisper-timestamped lintoai2023whispertimestamped takes audio $x_{\text{window}}$ as input and conducts automatic speech recognition with word-level timestamps; ii) GPT-4o achiam2023gpt is prompted in Chain-of-Thought wei2022chain to decompose keystep detection task into two sub-tasks ge-etal-2024-dkec: token classification and phrase span detection and ask LLMs to think step by step. The input for GPT-4o is the transcripts generated by Whisper-timestamped, and the output is the phrases from the transcripts, each labeled with keystep type. The detailed prompt is shown in Figure 34. iii) Keystep Boundary Detection further applies keystep span detection by grouping phrases with the same keystep label together and updating the start and end timestamp for each grouped keystep. The predictions for all segments $x_{\text{window}}$ are concatenated to form a complete trial, which is then passed to the final Keystep Boundary Detection module.
@@ D.3.1 Egocentric Setting
Supervised depth estimation. For this approach, we employ a supervised depth estimation model to estimate the depth of RGB frames (see Figure 36). Specifically, we utilize a monocular depth estimation method lasinger2019towards to obtain a relative depth map, which we scale to absolute depth using a reference object with known real-world depth. Further, we evaluate the large and small variants of the MiDas model introduced in lasinger2019towards.
@@
Rate (cpm) Depth (mm) F1 Acc F1 Acc
Egocentric (RGB) Heuristic (SP) 28.56 25.78 0.47 0.52 0.77 0.70
IMU (Acc) Supervised + SP 14.89 12.13 0.51 0.64 0.83 0.77
Fusion (Supervised) Supervised + SP 15.66 11.78 0.52 0.62 0.83 0.77
@@ D.3.2 Smartwatch IMU-Based Setting
Figures 37 and 38 present the average errors in compression rate and depth estimation for different modalities, grouped by participant background. In contrast to egocentric video based estimation, smartwatch IMU based method consistently achieve low errors for both EMS responders and general public participants. This improved performance stems from the direct and robust measurement of wrist motion by the IMU, which is largely unaffected by the occlusions, field-of-view limitations, and other challenges that hinder egocentric video based approaches. This is further evident from the substantial reduction in compression rate error for EMS responders when using IMU data compared to egocentric video. Following a similar trend, the smartwatch IMU based method achieves the highest feedback accuracy for compression rate at 0.64, and for compression depth at 0.77, outperforming other modalities in both metrics.
@@ D.3.3 Multimodal Fusion
Figure 39: Multimodal fusion approach for CPR quality estimation.
@@
These quantitative results are further supported by the box plots in Figures 37 and 38, which illustrate deviations in predicted CPR compression rate and depth across modalities and participant backgrounds. For both metrics, fusion methods exhibit consistently lower deviations compared to unimodal egocentric video, especially for EMS responders, whose results with ego video alone show large variability and frequent extreme errors. The addition of IMU data stabilizes predictions, reducing errors caused by wrist occlusions, camera motion, and confusion with secondary responders in multi-responder scenes.